A 8.50 g sample of copper is contaminated with 0.88 g of zinc. Suppose an atomic mass measurement was performed on this sample.What would be the measured atomic mass?
Answer #1

atomic mass of copper = 63.546 amu

atomic mass of zinc = 65.38 amu

total sample = 8.50 gm

zinc = 0.88gm

copper = 8.50 - 0.88 = 7.62 gm

Calculate % composition

% of copper = 7.62 \times100/8.50 = 89.64 %

% of zinc = 100 - 89.64 = 10.36 %

Avarage atomic mass of samle =( (atomic mass of copper \times % copper ) +(atomic mass of zinc \times % of zinc) )/100

( (63.546\times89.64) +(65.38\times10.36) )/100

(5696.26 +677.33)/100 = 63.7359

measured atomic mass of sample = 63.7359